# Graph Transformers without Positional Encodings

### v2

@@ Abstract
Recently, Transformers for graph representation learning have become increasingly popular, achieving state-of-the-art performance on a wide-variety of datasets, either alone or in combination with message-passing graph neural networks (MP-GNNs). Infusing graph inductive-biases in the innately structure-agnostic transformer architecture in the form of structural or positional encodings (PEs) is key to achieving these impressive results. However, designing such encodings is tricky and disparate attempts have been made to engineer such encodings including Laplacian eigenvectors, relative random-walk probabilities (RRWP), spatial encodings, centrality encodings, edge encodings etc. In this work, we argue that such encodings may not be required at all, provided the attention mechanism itself incorporates information about the graph structure. We introduce Eigenformer, a Graph Transformer employing a novel spectrum-aware attention mechanism cognizant of the Laplacian spectrum of the graph, and empirically show that it achieves performance comparable to SOTA Graph Transformers on a number of standard GNN benchmark datasets, even surpassing the SOTA on some datasets. The simpler attention mechanism also allows us to train wider and deeper models for a given parameter budget.
@@ 1 Introduction
Learning from graph data is useful in a variety of domains due to the ubiquitous presence of graphs around us, ranging from molecules to social networks and transportation networks. It may well be argued that "graphs are the main modality of data we receive from nature" [Veličković, 2023]. In general machine learning methods designed to learn patterns in graph data fall into two paradigms: message-passing graph neural networks (MP-GNNs), where vector messages are exchanged between neighboring nodes and updated using neural networks [Bruna et al., 2014, Gilmer et al., 2017], and more recently Transformers [Vaswani et al., 2017] –originally designed for Natural Language Processing (NLP) tasks– for graph data, which feature long-range connections between nodes for faster information exchange albeit with a loss of graph inductive biases [Dwivedi and Bresson, 2020, Kreuzer et al., 2021, Chen et al., 2022a, Ying et al., 2021, Rampášek et al., 2022, Hussain et al., 2022, Zhang et al., 2023].
@@
In many ways, MP-GNNs and Graph Transformers have complementary strengths and weaknesses. While MP-GNNs have strong graph inductive biases due to sharing of messages within local neighborhoods, they suffer from issues such as over-squashing [Alon and Yahav, 2021, Topping et al., 2022], over-smoothing [Li et al., 2018, Oono and Suzuki, 2020] and expressivity concerns (upper bounded by the 1-Weisfeiler-Lehman isomorphism test) [Xu et al., 2019, Loukas, 2020, Morris et al., 2019]. On the other hand, Graph Transformers do not suffer from these issues, at the expense of lacking graph inductive biases (forgoing locality with usually all-to-all connectivity between nodes), possibility of over-fitting and quadratic computational and memory complexity. Due to their complementary benefits, researchers have (successfully) tried to merge the two paradigms together into one architecture with GraphGPS in Rampášek et al. [2022].
@@
A common weakness in both methods however, is the lack of positional information of nodes, which has been shown to be crucial both to improve expressivity in MP-GNNs (as compared to the 1-WL test in Xu et al. [2019]) and to allow Transformers to be used with small-scale graph data. Nodes in a graph innately lack a canonical order, and superficially imposing an ordering leads to exponential growth in possible such orderings, leading to learning difficulties. To tackle this, PEs based on Laplacian eigenvectors were proposed in Dwivedi and Bresson [2020]. Due to sign ambiguities of eigenvectors and eigenvalue multiplicities, sign and basis invariant/equivariant methods such as SignNet and BasisNet in Lim et al. [2023] were proposed. In other works [Dwivedi et al., 2022a, Ma et al., 2023], k-dimensional (learnable) random-walk probabilities from a node to itself are used instead of Laplacian eigenvectors to avoid the aforementioned invariance issues.
@@
Contributions:   In light of the above background, we ask the question whether such encodings are indeed necessary at all. If there is no canonical node-ordering inherent to a graph, perhaps no such ordering is required to effectively learn patterns in graph data. Thus, in this work, we seek to circumvent the issues related to designing node/edge PEs by modifying the way node feature information is exchanged in a Graph Transformer. Specifically:
@@
  • Instead of adding positional information to node features and letting the attention mechanism figure out the strength of connection between a pair of nodes, we factorize the attention matrix in terms of fixed spectral similarities and learned frequency importances. We posit such a factorization encodes important graph inductive biases from the frequency domain into the attention mechanism, while being flexible enough to yield impressive performance across domains and objectives.
@@
  • We utilize valuable information from both the eigenvectors and the eigenvalues as advocated in Kreuzer et al. [2021], forgoing the assumption that higher frequencies are less important in determining node similarity.
@@
  • The novel attention mechanism with built-in inductive biases allows us to train deeper models for a given parameter budget due to simpler Transformer layers. Numerical experiments demonstrate the strength of our method, achieving near-SOTA performances on many common GNN benchmarks.
2 Related Work
Graph connectivity and graph topology information can be very important for a GNN architecture, and following the introduction of Graph Transformers in Li et al. [2019], there have been many attempts to incorporate graph structural information in Transformers.
Graph inductive biases in TransformersDwivedi and Bresson [2020] perform attention only over the 1-hop neighborhood of nodes in addition to the use of Laplacian eigenvectors as node PEs. With Graph-BERT, Zhang et al. [2020] propose to train Transformers using sampled linkless (no edges) subgraphs within their local contexts, designed with an emphasis on parallelization and mainly as a pretraining technique. They use a number of PEs like the WL Absolute Role Embedding, Intimacy based Relative Positional Embedding and Hop based Relative Distance Embedding, in order to retain the original graph structural information in the linkless subgraphs.
Kreuzer et al. [2021] use a Transformer encoder to create learnable PEs from a combination of eigenvalues and eigenvectors which are then fed to the main Transformer network. Dwivedi et al. [2022a] decouple the node features and the node PEs with a separate propagation step for the PEs. They consider two different initializations for the PEs: Laplacian PE (LapPE) and Random Walk PE (RWPE), and show that RWPE outperforms LapPE, indicating that learning the sign invariance ambiguity in LapPE is more difficult than not exactly having unique node representation for each node. Ying et al. [2021] propose the use of a fixed centrality encoding to inject degree information in their Graphormer.
With GraphiT, Mialon et al. [2021] enumerate, encode and aggregate local sub-structures such as paths of length-k using Graph convolutional kernel networks (GCKN) [Chen et al., 2020] to create node PEs. Ma et al. [2023] propose GRIT, again using PEs initialized with random walk probabilities but make them flexible and learnable by composing with an MLP, and provide theoretical proof of the expressivity of such a composition. Further, they inject degree information using an adaptive degree-scaler [Corso et al., 2020] in every layer, noting that attention mechanisms are invariant to node-degrees.
Structure-aware attention mechanisms   More similar to our work, there has also been research on creating structure-aware attention mechanisms in order to infuse node structural and neighborhood similarity information. Mialon et al. [2021] view self-attention as kernel smoothing and bias the attention scores using diffusion or k-step random-walk kernels. Chen et al. [2022b] argue that such a modification is deficient in filtering out structurally dissimilar nodes, and consider a more generalized kernel accounting for local substructures like k-subtrees and k-subgraphs, around each node. In a similar approach, G et al. [2023] aggregate node feature information from k-subtrees but augment their transformer with a global node and train positional encodings using the Neighbourhood Contrastive Loss (NCE). Ying et al. [2021] have proposed to inject spatial and edge information by adding learnable encodings to the attention scores before normalization, while [Menegaux et al., 2023] augment that idea further by learning a different attention score per hidden dimension in their so-called Chromatic Self-Attention. Learned edge encodings are also incorporated in Hussain et al. [2022] both as a bias to the attention score before normalization and as a sigmoid-gate after normalization.
In contrast to previous work, to the best of our knowledge, there have been no attempts to factorize the attention matrix into spectral similarities and frequency importances, which we present in detail in the next section.
3 Methodology and Theory
In this section, we introduce our novel spectrum-aware attention (SAA) mechanism and its use in the proposed Eigenformer architecture. We first present the motivation behind the architecture by reproducing the meaning of the graph Laplacian and its spectrum.
@@ 3.1 Graph Laplacian and its spectrum
where $A$ is the adjacency matrix and $D$ is the degree matrix.
By the Rayleigh-Ritz theorem, the eigenvectors $u_{i},i=1,...,|\mathcal{V}|$ of the Laplacian solve the problems:
$argmin_{u_{i}\in\mathbb{R}^{|\mathcal{V}|}:u_{i}\perp u_{j}\forall j<i}\frac{u_{i}^{T}Lu_{i}}{u_{i}^{T}u_{i}}$ (2)
and the values attained are the eigenvalues $\lambda_{i},i=1,...,|\mathcal{V}|$.
By definition 1, the following is also true for a unit-norm eigenvector $u_{i}$ and its corresponding eigenvalue $\lambda_{i}$:
@@
If the eigenvectors are interpreted as "signals" over the graph, taking real scalar values at nodes, then the eigenvalues act as a measure of the "smoothness" of such signals over the graph, measuring how much the signal values differ between endpoints of edges. The smallest eigenvalue is always 0, assigning the same scalar value $u_{i}[\cdot]$ to all nodes in the same connected component of the graph.
As discussed in previous works (eg. Kreuzer et al. [2021]), the Laplacian spectrum equates to sine functions over graphs and encodes important information about relative distances between nodes. Nodes close to each other tend to have similar values under a given eigenvector. As such, the spectrum –in its interpretation as frequencies of resonance of the graph– discriminates between different graph structures and sub-structures. Figure 1 shows substructures revealed by eigenvectors for an example graph from the ZINC dataset containing 33 nodes. Color-coded nodes indicate the value $u_{i}[\cdot]$ for different frequencies $\lambda_{i}$ where $i\in[0,32]$.
(a) $u_{1}$
(b) $u_{4}$
(c) $u_{12}$
Figure 1: ZINC dataset: Graph substructures revealed by the eigenvectors
There is also a normalized variant of the graph Laplacian, $L_{sym}=D^{-\frac{1}{2}}LD^{-\frac{1}{2}}$ that admits similar properties and interpretations of the spectrum as that of $L$, but has better numerical stability due to bounded spectrum: all eigenvalues of $L_{sym}$ lie between 00 and $2$. Moreover, the normalized Laplacian counters the effect of "heavy" nodes (nodes with high degrees) in influencing information propagation through the graph. We propose to use the spectrum of $L_{sym}$ to define a new spectrum-aware attention mechanism as outlined in the next section.
@@ 3.2 Attention using the Laplacian spectrum
We define the attention weight between two nodes $i$ and $j$ through the following equations:
@@
where $u_{k}$ is the eigenvector corresponding to the eigenvalue $\lambda_{k}$ of $L_{sym}$, $\sigma_{k}$ is assumed to measure the similarity between nodes "under" the frequency $\lambda_{k}$, $\phi_{1}$ and $\phi_{2}$ are arbitrary functions and $\alpha$ is the final attention weight.
The matrices $\sigma_{k}[\cdot,\cdot]$ for $k=1...|\mathcal{V}|$ can be thought of as defining $|\mathcal{V}|$eigen-similarity" matrices measuring the strength of connection between pairs of nodes under different frequencies. Nodes simultaneously having high values and same sign of $u_{k}[\cdot]$ are deemed to be more correlated under that frequency. The justification for this choice of $\sigma_{k}$ is presented in Proposition 1 below (proof in Appendix A).
@@
For any $n\in\mathbb{N}$, consider the symmetric adjacency matrix $A$ drawn from the set of adjacency matrices of n-node undirected graphs, $\mathbb{G}_{n}\subset\{0,1\}^{n\times n}$. Further, let $L_{sym}=I-D^{-\frac{1}{2}}AD^{-\frac{1}{2}}=I-A_{sym}$ be the normalized graph Laplacian of the graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with adjacency matrix $A$, eigenvalues $\lambda_{k}$ and eigenvectors $u_{k}$ for $k\in[1,|\mathcal{V}|]$ with $|\mathcal{V}|=n$. Then, we have the following approximations upto an arbitrary small error $\epsilon$:
@@
    $\phi_{1}\big{(}\sum_{k=1}^{|\mathcal{V}|}\sigma_{k}[i,j]\cdot\phi_{2}(\lambda_{k})\big{)}\approx\sum_{k=0}^{m}\theta_{k}A_{sym}^{k}[i,j]$ for $m\in\mathbb{Z}$
@@
    $\phi_{1}\big{(}\sum_{k=1}^{|\mathcal{V}|}\sigma_{k}[i,j]\cdot\phi_{2}(\lambda_{k})\big{)}\approx SPD[i,j]$
for suitable functions $\phi_{1}$ and $\phi_{2}$, where $SPD[i,j]$ is the shortest path distance between nodes $i$ and $j$.
Proposition 1 states that the attention mechanism defined in 5 can capture (i) shortest path distances between all pairs of nodes and (ii) all weighted combinations of powers of the normalized adjacency matrix $A_{sym}$ allowing us to perform aggregations over any $k-$hop neighborhood where $k\in\mathbb{N}$. The function $\phi_{2}$ quantifies the importance of the frequency $\lambda_{k}$ in the sum in 5, deciding how to mix the $|\mathcal{V}|$ eigen-similarity matrices to obtain node connection strengths that are useful for the graph learning task at hand. Traditional distance metrics like the biharmonic distance [Lipman et al., 2010] or the diffusion distance [Coifman and Lafon, 2006, Bronstein et al., 2017], weigh smaller frequencies more heavily, assuming that higher frequencies play a smaller role in the measurement of relative distance between nodes. We make no such assumption, rather allowing the dependence on the frequency to be learnt from data by parameterizing the function $\phi_{2}$ appropriately.
Note that calculation of the similarities in 4 is possible offline with just the graph adjacency matrix $A$ whereas 5 requires us to choose the form of $\phi_{1}$ and $\phi_{2}$. Thus, the only learnable components in our proposed method are the functions $\phi_{1}$ and $\phi_{2}$, leading to a drastic simplification of the attention mechanism.

### v3

@@ Abstract
Recently, Transformers for graph representation learning have become increasingly popular, achieving state-of-the-art performance on a wide-variety of graph datasets, either alone or in combination with message-passing graph neural networks (MP-GNNs). Infusing graph inductive-biases in the innately structure-agnostic transformer architecture in the form of structural or positional encodings (PEs) is key to achieving these impressive results. However, designing such encodings is tricky and disparate attempts have been made to engineer such encodings including Laplacian eigenvectors, relative random-walk probabilities (RRWP), spatial encodings, centrality encodings, edge encodings etc. In this work, we argue that such encodings may not be required at all, provided the attention mechanism itself incorporates information about the graph structure. We introduce Eigenformer, a Graph Transformer employing a novel spectrum-aware attention mechanism cognizant of the Laplacian spectrum of the graph, and empirically show that it achieves performance competetive with SOTA Graph Transformers on a number of standard GNN benchmarks. Additionally, we theoretically prove that Eigenformer can express various graph structural connectivity matrices, which is particularly essential when learning over smaller graphs.
@@ 1 Introduction
Learning useful representations from graph data is important in a variety of domains due to the ubiquitous presence of graphs around us, ranging from molecules to social and transportation networks. It may well be argued that "graphs are the main modality of data we receive from nature" [Veličković, 2023]. In general, machine learning methods designed to learn patterns from graph data fall into two paradigms: message-passing graph neural networks (MP-GNNs), where vector messages are exchanged between neighboring nodes and updated using neural networks [Bruna et al., 2014, Gilmer et al., 2017], and more recently Transformers [Vaswani et al., 2017] –originally designed for Natural Language Processing (NLP) tasks– for graph data, which feature long-range connections between nodes for faster information exchange albeit with a loss of graph inductive biases [Dwivedi and Bresson, 2020, Kreuzer et al., 2021, Chen et al., 2022a, Ying et al., 2021, Rampášek et al., 2022, Hussain et al., 2022, Zhang et al., 2023].
@@
In many ways, MP-GNNs and Graph Transformers have complementary strengths and weaknesses. While MP-GNNs have strong graph inductive biases due to sharing of messages within local neighborhoods, they suffer from issues such as over-squashing [Alon and Yahav, 2021, Topping et al., 2022], over-smoothing [Li et al., 2018, Oono and Suzuki, 2020] and expressivity concerns (upper bounded by the 1-Weisfeiler-Lehman isomorphism test) [Xu et al., 2019, Loukas, 2020, Morris et al., 2019]. On the other hand, Graph Transformers do not suffer from these issues, at the expense of losing graph connectivity information (forgoing locality in favor of all-to-all connectivity between nodes), possibility of over-fitting and quadratic computational and memory complexity. Due to their complementary benefits, researchers have (successfully) tried to merge the two paradigms together into one architecture with GraphGPS in Rampášek et al. [2022].
@@
A common weakness in both methods, however, is the lack of node positional information, which has been shown to be crucial both to improve expressivity in MP-GNNs (as compared to the 1-WL test in Xu et al. [2019]), and to allow Transformers to be used with datasets with smaller graphs. Nodes in a graph lack a canonical ordering, and superficially imposing an ordering leads to learning difficulties due to the exponential growth in such possible orderings. To tackle this, positional encodings (PEs) based on Laplacian eigenvectors were proposed in Dwivedi and Bresson [2020]. Due to sign ambiguities of eigenvectors and eigenvalue multiplicities, sign and basis invariant/equivariant methods such as SignNet and BasisNet in Lim et al. [2023] were proposed. In other works [Dwivedi et al., 2022a, Ma et al., 2023], $k$-dimensional (learnable) random-walk probabilities from a node to itself are used instead of Laplacian eigenvectors to avoid the aforementioned invariance issues.
@@
Main contributions: In light of the above background, we ask the question whether such encodings are indeed necessary at all. If there is no canonical node-ordering inherent to a graph, perhaps no such ordering is required to effectively learn patterns from graph data. Thus, in this work, we seek to circumvent the issues related to designing node/edge PEs by modifying the way node feature information is exchanged in a Graph Transformer. Specifically:
@@
  • Instead of adding positional information to node features and letting the attention mechanism figure out the strength of connection between a pair of nodes, we factorize the attention matrix in terms of fixed node-pair "potentials" and learned frequency "importances". We posit such a factorization encodes important graph inductive biases from the frequency domain into the attention mechanism while being flexible enough to yield impressive performance across domains and objectives.
@@
  • We utilize valuable information from both the eigenvectors and the eigenvalues as advocated in Kreuzer et al. [2021], forgoing the assumption that higher frequencies are less important in determining node-pair interactions.
@@
  • The novel attention mechanism with built-in inductive biases allows us to train wider/deeper models for a given parameter budget due to simpler Transformer layers. Numerical experiments demonstrate the strength of our method, achieving near-SOTA performances on many common GNN benchmarks.
2 Theoretical Motivations
We first present the motivation behind the architecture of Eigenformer by briefly reviewing the semantics of the graph Laplacian, its spectrum and spectral convolution.
@@ 2.1 Graph Laplacian and its spectrum
where $A$ is the adjacency matrix and $D$ is the degree matrix. By definition 1, the following is true for a unit-norm eigenvector $u_{i}$ and its corresponding eigenvalue $\lambda_{i}$:
@@
If the eigenvectors are interpreted as "signals" over the graph, taking real scalar values at nodes, then the eigenvalues act as estimates of the "smoothness" of such signals, measuring how much the signal values differ between endpoints of edges. The smallest eigenvalue is always 0, assigning the same scalar value $u_{i}[\cdot]$ to all nodes in the same connected component of the graph.
Thus, eigenvectors of the Laplacian encode important information about node similarities (nodes resonant under a frequency tend to have similar values w.r.t its eigenvector). As such, the spectrum –in its interpretation as frequencies of resonance of the graph– discriminates between different graph structures and substructures. Figure 1 shows substructures revealed by eigenvectors for an example molecular graph from the PCQM4Mv2 dataset [Hu et al., ] containing 15 nodes. Color-codes indicate the value $u_{i}[\cdot]$ for different frequencies $\lambda_{i}$ where $i\in\{0,...,15\}$. Lower frequencies highlight larger substructures while higher frequencies emphasize more local interactions.
Figure 1: Example molecule from PCQM4Mv2 dataset: Substructures are revealed by eigenvectors
2.2 Spectral convolution
Since the (sinusoidal) eigenfunctions $f(t)=e^{2\pi ist}$, of the general Laplace operator $\Delta$, correspond to the modes of the frequency domain of the Fourier transform, the eigenvectors of the graph Laplacian $L$ can be equated to sine functions over graphs. This connection allows us to generalize the Fourier transform to arbitrary graphs.
For a signal $f\in\mathbb{R}^{|\mathcal{V}|}$ on an undirected graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with Laplacian $L=U\Lambda U^{T}$, where $U$ is the orthonormal matrix of eigenvectors and $\Lambda$ is the diagonal matrix of eigenvalues, we can compute its Fourier transform as $s=U^{T}f$, and its inverse as $f=Us$. Thus, its graph convolution $\star_{\mathcal{G}}$, with a filter $h$ can be computed as:
$f\star_{\mathcal{G}}h=U(U^{T}f\odot U^{T}h)=(U\text{diag}(w_{h})U^{T})f$ (3)
where the last equality is true when the filter $h$ can be learnt in a non-parametric way by representing and optimizing directly in the spectral domain: $w_{h}=U^{T}h$. Such a filter, however, may not satisfy sound graph convolution properties like locality and translational equivariance, and it is useful to parameterize $w_{h}$ in terms of eigenvalues $\Lambda$ of the Laplacian, for example, as a polynomial $p(\Lambda)$, which ensures commutativity with the Laplacian, restoring the required properties:
$f\star_{\mathcal{G}}h=(Up(\Lambda)U^{T})f=p(L)f$ (4)
Thus, a spectral convolution on the graph can be defined in terms of a polynomial of the Laplacian $L$. The degree $k$, of the polynomial determines the $k$-hop neighborhood over which the convolution is performed.
3 Model Architecture
In this section, we introduce our novel spectrum-aware attention (SAA) mechanism that builds on the spectral convolution theory discussed in 2.2 and demonstrate its use in the proposed Eigenformer architecture.
Note: We use the normalized variant of the graph Laplacian, $L_{norm}=D^{-\frac{1}{2}}LD^{-\frac{1}{2}}$ that admits similar properties and interpretations of the spectrum as that of $L$, but has better numerical stability due to a bounded spectrum: all eigenvalues of $L_{norm}$ lie between 00 and $2$. Moreover, the normalized Laplacian counters the effect of "heavy" nodes (nodes with high degrees) in influencing information propagation through the graph. Also, for directed graphs, we first make the edges undirected before calculating $L_{norm}$ so that its spectrum remains real.
@@ 3.1 Attention using the Laplacian spectrum
For a graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ we define the attention weight between two nodes $i$ and $j$ in $\mathcal{G}$, through the following equations:
@@
where $u_{k}\in\mathbb{R}^{|\mathcal{V}|}$ is the eigenvector corresponding to the eigenvalue $\lambda_{k}\in\mathbb{R}$ of $L_{norm}$, $\sigma_{k}\in\mathbb{R}$ can loosely be interpreted as the "potential" between the nodes under the frequency $\lambda_{k}$ (see Green’s function from electrostatic theory, Chung and Yau [2000]), $\phi_{1}:\mathbb{R}\rightarrow\mathbb{R}$ and $\phi_{2}:\mathbb{R}\rightarrow\mathbb{R}$ are arbitrary real functions and $\alpha\in\mathbb{R}$ is the final attention weight. The connection to spectral convolution can be inferred by noting that $\alpha=softmax[\phi_{1}(U\phi_{2}(\Lambda)U^{T})]$ where $U=[u_{1},...,u_{|\mathcal{V}|}]$.
Propositions 1 and 2 below justify the choice of $\sigma$ and $\alpha$ rigorously.
@@
For any $n\in\mathbb{N}$, consider the adjacency matrix $A$ drawn from the set of adjacency matrices of n-node undirected graphs, $\mathbb{G}_{n}\subset\{0,1\}^{n\times n}$. Further, let $L_{norm}=I-D^{-\frac{1}{2}}AD^{-\frac{1}{2}}=I-A_{norm}$ be the normalized graph Laplacian of the graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$ with $|\mathcal{V}|=n$, adjacency matrix $A$, eigenvalues $\lambda_{k}$ and eigenvectors $u_{k}$, for $k\in\{1,...,n\}$. Then, we have the following approximations up to an arbitrary small error, $\epsilon$:
@@
    $\phi_{1}\big{(}\sum_{k=1}^{|\mathcal{V}|}\sigma_{k}[i,j]\cdot\phi_{2}(\lambda_{k})\big{)}\approx\sum_{k=0}^{m}\theta_{k}A_{norm}^{k}[i,j]$ for $m\in\mathbb{Z}$
@@
    $\phi_{1}\big{(}\sum_{k=1}^{|\mathcal{V}|}\sigma_{k}[i,j]\cdot\phi_{2}(\lambda_{k})\big{)}\approx f(SPD[i,j])$
for suitable functions $\phi_{1}$ and $\phi_{2}$, where $SPD[i,j]$ is the shortest path distance between nodes $i$ and $j$ and $f$ is any continuous function.
Proof: The proof of (1.) utilizes the diagonalizability of the matrices $A_{norm}$ and $L_{norm}$, allowing us to derive the analytic form of $\phi_{1}$ and $\phi_{2}$. For (2.), we show that $\phi_{2}$ can be chosen to yield a valid (possibly many-to-one) function that maps $\sum_{k=1}^{|\mathcal{V}|}\sigma_{k}[i,j]\cdot\phi_{2}(\lambda_{k})$ to $f(SPD[i,j])$ for all node-pairs $(i,j)$ in a graph $\mathcal{G}$ in the family of graphs indexed by $\mathbb{G}_{n}$, and we define $\phi_{1}$ to be that piecewise-linear map. The detailed analysis can be found in Appendix A.1).
Proposition 1 states that the attention mechanism defined in 6 can capture (i) shortest path distances between all pairs of nodes and (ii) all weighted combinations of powers of the normalized adjacency matrix $A_{norm}$ allowing us to perform aggregations over any $k-$hop neighborhood where $k\in\mathbb{Z}$. The function $\phi_{2}$ quantifies the "importance" of the frequency $\lambda_{k}$ in the sum in 6. Traditional distance metrics like the biharmonic distance [Lipman et al., 2010] or the diffusion distance [Coifman and Lafon, 2006, Bronstein et al., 2017], weigh smaller frequencies more heavily, assuming that higher frequencies play a smaller role in the measurement of relative distance between nodes. We make no such assumption, rather allowing the dependence on the frequency to be learnt from data by parameterizing the function $\phi_{2}$ appropriately.
We now show that the proposed attention mechanism does not suffer from eigenvector sign/basis-invariance issues, making prior interventions like SignNet/BasisNet [Lim et al., 2023] redundant.
Proposition 2
The attention matrix $\alpha$ computed using equations 5 and 6, is invariant to the choice of the sign of the eigenvectors $u_{k}$, and in the general case, to the choice of basis of the eigenspace corresponding to the eigenvalue $\lambda_{k}$ having geometric multiplicity > 1.
Proof: The sign-invariance can be easily inferred by noting that eigenvectors $u_{k}$ appear in the computation of $\alpha$ as multiplied pairs of terms (5). The invariance to basis is due to the fact that in 6, the eigenvectors $U_{k}=[u_{k_{1}},...,u_{k_{d_{i}}}]\in\mathbb{R}^{n\times d_{i}}$ corresponding to the same eigenvalue $\lambda_{k}$ occur together as terms of the form $h(U_{k}U_{k}^{T})$, for some continuous function $h$. Since any other orthonormal basis is of the form $V_{k}=U_{k}Q$ for some orthogonal $Q\in O(d_{i})\subseteq\mathbb{R}^{d_{i}\times d_{i}}$, we have $h(V_{k}V_{k}^{T})=h(U_{k}U_{k}^{T})$, implying that the choice of basis does not affect the computation in 6.